Mr. Nazri saves RM500 at the beginning of each year. He receives an interest of 3% each year. Find the total amount of savings including the interest after 20 years


1
Expert's answer
2021-09-28T10:42:09-0400
"F=P(\dfrac{(1+r)^n-1}{r})"

"F=500(\dfrac{(1+0.03)^{20}-1}{0.03})=13435.19"

The total amount of savings including the interest after 20 years is "RM13435.19"